# Flag complex face structures and decompositions

### v1

@@ 2.1 Background related to cell and cubical complexes
We give some background on objects that we will be studying. Recall that the $w=2\alpha+1$ substitution was motivated by trying to use $f_{S}(x)=F_{S}(2x+1)$ for a simplicial complex $S$ to show that the input for the Tchebyshev transformation (for polynomials) in $T(P_{\Delta}(w))=g_{\Delta}(2w)=(2w+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2w+2}\right)$ is the $F$-vector of some simplicial complex. However, the indices from Theorem 2.22 are flipped in that we have the index $\frac{d}{2}-k-1$ term for the coefficient of $\alpha^{k}$ instead of something of index $k-1$.
@@
A cubical complex that is closely connected to $F$-polynomials we are considering the mirroring/Danzer complex $M\Delta$ of a cell complex $\Delta$, which is a cubical complex whose vertices have links isomorphic to $\Delta$.
@@
    By cell complex, we mean a polyhedral complex or the (polar) dual of one (analogous to polar duals of polytopes).
@@
    We will take the dual $A^{*}$ of a cell complex $A$ of dimension $D-1$ to mean the cell complex where a face of dimension $j$ corresponds to the collection of facets of $A$ that contain a particular $(D-j-1)$-dimensional face of $A$. This is an analogue of the polar dual of a polytope.
We can also give a generalization of polar duality for multicomplexes.
Definition 2.2.
  1. 1.
    (Definition 2.4 on p. 25 – 26 of [6]) A multicomplex is a collection of monomials $\mathcal{M}$ such that $m|m^{\prime}\in\mathcal{M}$ implies that $m\in\mathcal{M}$ for all monomials $m^{\prime}\in\mathcal{M}$.
    Let $\mathcal{M}^{r}$ denote the set of all monomials in $\mathcal{M}$ of degree $r$. The $f$-vector of a multiplex $\mathcal{M}$ is the sequence $(f_{-1},f_{0},f_{1},\ldots)$ where $f_{j-1}(\mathcal{M})=|\mathcal{M}^{j}|$ for $j=0,1,2,\ldots$. Note that this uses a different indexing convention from [6].
  2. 2.
    The dual $\mathcal{M}^{*}$ of a degree $D$ multicomplex $\mathcal{M}$ is the complex where the 0-dimensional faces/vertices come from maximal degree monomials and the $(k-1)$-dimensional faces of $\mathcal{M}^{*}$ come from quotients of degree $D$ terms of $\mathcal{M}$ by degree $D-k$ terms of $\mathcal{M}$ In other words, the“vertices” contained in a face $R^{*}\in\mathcal{M}$ of degree $k$ corresponding to a monomial $R\in\mathcal{M}$ come from to monomials of maximal degree in $\mathcal{M}$ that are divisible by $R$.
@@
partially ordered by $0>1$ and $0>-1$ extended componentwise. Here $I$ is the poset $\{0,1,-1\}$ with ordering $1<0$ and $-1<0$. Note that $\dim MT=\dim T+1$.
@@
The way we will think about the notation above is that placing $1$ or $-1$ in the $i^{\text{th}}$ coordinate refers to intersecting with the hyperplane $x_{i}=1$ or $x_{i}=-1$ defining a facet of the $n$-cube $[-1,1]^{n}$ and that the coordinates labeled with a “0” are “free” in the sense that they are used to span a face of the cubical complex $MT$. In Rowlands’ thesis (see p. 4 of [28]), the “free” coordinates are labeled with a “*”, which may better convey this. As a consequence, the faces of $MT$ with 00’s in the same coordinates (i.e. coming from the same face of $T$) are parallel to each other.
@@
    Suppose that the faces $G_{1},\ldots,G_{m}\in MT$ have a nonempty intersection. For each $1\leq i\leq m$, let $F_{i}=\{(|a_{1}|,\ldots,|a_{n}|):(a_{1},\ldots,a_{n})\in G_{i}\}$ (i.e. subset labeled by 0’s corresponds to $F_{i}$). Then, we have that $G_{1}\cap\cdots\cap G_{m}$ has 0’s in the same slots as $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the “common values” from Part 1 for coordinates corresponding to elements of the complement $[n]\setminus(F_{1}\cap\cdots\cap F_{m})$. In particular, this means that $G_{1}\cap\cdots\cap G_{m}\in M\Delta$ since the subset of $[n]$ labeled with 0’s is the face $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the rest of the coordinates are labeled with ±$\pm 1$.
@@
    • $\operatorname{st}_{M\Delta}(G)\cong\operatorname{st}_{\Delta}(|G|)$, where $|G|$ denotes the face of $\Delta$ obtained after taking the absolute value of each coordinate (i.e. setting the $-1$ coordinates of $G$ equal to 1).
    • The nonzero coordinates of elements of $\operatorname{st}_{M\Delta}(G)$ are determined by those of $G$ (which are indexed by $[n]\setminus|G|$).
Returning to the original question of the interpretation of $P_{\Delta}(2\alpha+1)$, we use the following result connecting faces of a cell complex with its mirroring/Danzer complex.
@@
Suppose that we include the empty set in the face poset of a cell complex $T$ and that it is not included in the face poset of $MT$. If every $(k-1)$-dimensional cell of $T$ has $k$ vertices, we have that
@@
2.2 Colorings, Tchebyshev subdivisions, and intervals
Now suppose that $\gamma(\Delta)=f(T)$ for some simplicial complex $T$ of dimension $\frac{d}{2}-1$. By Theorem 0.3, we have that
@@ 2.2 Colorings, Tchebyshev subdivisions, and intervals
The first cubical barycentric subdivision of a simplicial complex $S$ of dimension $D-1$ is the cubical complex $\mathcal{C}(S)$ defined as follows:
@@
    $[\sigma,\tau]\coloneq\{\lambda\in S:\sigma\subset\lambda\subset\tau\}.$
@@
(Comparison with the formula for $[w^{k}]g(2w)$ given if $\gamma(\Delta)=f(T)$)
@@
    Comparing this to the formula for $[w^{k}]g(2w)$ obtained when $\Gamma(\Delta)=f(T)$, we see that it uses $\binom{j}{k}$ instead of $\binom{j+1}{k}$ and has an input that might not be a simplicial complex. Note that the $k$-dimensional faces of $MA$ come from $(k-1)$-dimensional faces of $A$ (see Definition 2.3).
@@
    The cubical barycentric subdivision assumes that both endpoints of the interval considered are contained in $S$. If we drop the condition that $S$ is a simplicial complex, we’re simply considering endpoints $[\sigma,\tau]$ where the upper endpoint is a $j$-dimensional face of $S$ and $\tau$ is some subset of $\sigma$ (which may or may not be a face of $S$).
@@
Given a poset $P$, its barycentric cover $KP$ is the poset with elements the order relations of $P$, partially ordered by inclusion (i.e. $(u\leq v)\leq(x\leq y)$ (see Definition 2.14) if and only if $x\leq u\leq v\leq y$). Note that $KP=K^{\text{op}}$ (see p. 298 of [5]) and $K(P\times Q)=K(P)\times K(Q)$.
@@
An interval $[u,v]$ in a partially ordered set $P$ is the set of all elements $w\in P$ satisfying $u\leq v\leq w$. For a finite partially ordered set $P$, we define the poset $I(P)$ of the nonempty intervals of $P$ as the set of all intervals $[u,v]\subset P$, ordered by inclusion.
@@
For each $u\in P$, we identify the vertex $[u,u]\in\Delta(I(P))$ with the vertex $u\in\Delta(P)$ and for each nonsingleton interval $[u,v]\in I(P)$, we identify the vertex $[u,v]\in\Delta(I(P))$ with the midpoint of the edge $\{[u,u],[v,v]\}$. We number the midpoints $[u_{1},v_{1}],[u_{2},v_{2}],\ldots$ in such a way that $i<j$ holds whenever the interval $[u_{i},v_{i}]$ contains the interval $[u_{j},v_{j}]$.
@@
Let $P$ be a graded poset and $\widehat{I}(P)$ be its poset of intervals. Then the order complex $\Delta(\widehat{I}(P)-\{\emptyset,[\widehat{0},\widehat{1}]\})$ is a Tchebyshev triangulation of the suspension of $\Delta(P-\{\widehat{0},\widehat{1}\})$.
@@
We take a page from p. 13 of Rowlands’ paper [29], which shows that $\widetilde{f}(\Box_{P},t)=\widetilde{f}(\Delta_{P},t+1)$. Note that any flag complex is the crossing complex $\Delta_{P}$ of a $PIP$ $P$ (Lemma 3.3 on p. 11 of [29]) and any rooted $CAT(0)$ cubical complex is of the form $\Box_{P}$ for some $P$ (Theorem 2.9 on p. 9 of [29]). The proof uses a bijection between dimension $i$ elements of $\Box_{P}$ and pairs $(A,M)$, where $A$ is the set of consistent antichains of $P$ and $M\subset A$ is a subset with $|M|=i$. The identity follows after making a change in the order of summation.
@@
$[w^{k}]\widetilde{f}\left(T(A),\frac{w-1}{2}\right)=\sum_{j=k}^{\frac{d}{2}}\widetilde{f}_{j}(T^{*})2^{j}\binom{j}{k},$
where $A$ is cell complex related to unused colors of faces in a balanced simplicial complex with an assigned coloring. Note that the last expression makes sense since a $(k-1)$-face of $T(A)$ has $k$ vertices (see Definition 2.26).
@@
Let $\Gamma$ be a $(D-1)$-dimensional balanced simplicial complex. Fix a $D$-coloring of the $1$-skeleton of $\Gamma$. We define the signed unused color complex $\mathcal{D}(\Gamma)$ of $\Gamma$ to be the cell complex below.
@@
  • The $(k-1)$-faces of the form $B\subset Q\subset C_{F}$, where $C_{F}$ is the set of colors in $[D]$ not used by $F$, $Q\subset C_{F}$ is a $k$-element subset, and $B\subset Q$ is an arbitrary subset. This is like considering “signed” $k$-element subsets of $C_{F}$. Note that $C_{F}\neq\emptyset$ since $Q$ is nonempty for the cases we will consider.
@@
Let $\Gamma$ be a balanced simplicial complex of dimension $D-1$. A collection of vertices $v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\in V(\mathcal{D}(\Gamma))$ with $\varepsilon_{i}=\pm 1$ forms a face of $\mathcal{D}(\Gamma)$ if and only if one of the following conditions (in the notation of Definition 2.19) holds:
@@
The second condition is nontrivial. For example, suppose that $\Gamma$ is the boundary of the octahedron/2-cross polytope (meaning $D=3$). This is a balanced simplicial complex. Color the vertices ±$\pm e_{i}$ with $i$ for $1\leq i\leq 3$ by $i$ to get a 3-coloring of its 1-skeleton. Consider the vertex $F_{1}=\{-e_{2}\}$ and the edge $F_{2}=\{e_{2},e_{3}\}$. Then, we have that $C_{F_{1}}=\{1,3\}$ and $C_{F_{2}}=\{1\}$. Since $F_{1}\cap F_{2}=\emptyset$ and $C_{F_{1}}\cup C_{F_{2}}=\{1,3\}$, the union of any two vertices of $\mathcal{D}(\Gamma)$ associated to these faces cannot form a face of $\mathcal{D}(\Gamma)$.
@@
Each vertex is of the form $B_{i}\subset\{q_{i}\}\subset C_{F_{i}}$, where $C_{F_{i}}$ is the set of unused colors of $F_{i}\in\Gamma$, $q_{i}\in C_{F_{i}}$, and $B_{i}=\{q_{i}\}$ or $\emptyset$. Using the inclusion relation in Definition 2.19, we have that the union $\{v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\}$ is given by $B_{1}\cup\cdots\cup B_{k}\subset\{q_{1},\ldots,q_{k}\}\subset(C_{1}\cup\cdots\cup C_{k})_{F_{1}\cap\cdots\cap F_{k}}$. If $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$, this forms a face of $\mathcal{D}(\Gamma)$ since $(P\cap Q)^{c}=P^{c}\cup Q^{c}$ for $P,Q\subset[D]$. Now suppose that $F_{1}\cap\cdots\cap F_{k}=\emptyset$. As mentioned in Definition 2.19, we have that $C_{\emptyset}=[D]$. This is only possible if $C_{F_{1}}\cup C_{F_{k}}=[D]$.
@@
$[\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}-\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)+1=f_{k-1}(\mathcal{D}(\Gamma))$
and
@@
for a cell complex $A$ of dimension $D-1$.
@@
Since $\Delta$ is a flag simplicial homology sphere, it is also a flag Cohen–Macaulay simplicial complex (e.g. see p. 21 of [4]). Then, a combining a result of Caviglia–Constantinescu–Varbaro (Corollary 2.3 on p. 474 of [8]) with that of Björner–Frankl–Stanley ($a=(1,\ldots,1)$ case of Theorem 1 on p. 23 and p. 33 of [6]) implies that $h(\Delta)$ is the $f$-vector of a balanced $\left(\frac{d}{2}-1\right)$-dimensional simplicial complex $\Gamma$.
@@
$P_{\Delta}(2\alpha+1)$ $=\sum_{i=0}^{\frac{d}{2}}f_{i-1}(\Gamma)(2\alpha+1)^{\frac{d}{2}-i}$
$=\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)(2\alpha+1)^{j}.$

### v2

@@ 1.2.1 Background related to cell and cubical complexes
We give some background on objects that we will be studying. Recall that the $w=2\alpha+1$ substitution was motivated by trying to use $f_{S}(x)=F_{S}(2x+1)$ for a simplicial complex $S$ to show that the input for the Tchebyshev transformation (for polynomials) in $T(P_{\Delta}(w))=g_{\Delta}(2w)=(2w+2)^{\frac{d}{2}}\gamma_{\Delta}\left(\frac{1}{2w+2}\right)$ is the $F$-vector of some simplicial complex. However, the indices from Theorem 1.25 are flipped in that we have the index $\frac{d}{2}-k-1$ term for the coefficient of $\alpha^{k}$ instead of something of index k1k-1.
@@
A cubical complex that is closely connected to $F$-polynomials we are considering the mirroring/Danzer/power complex $M\Delta$ of a cell complex $\Delta$, which is a cubical complex whose vertices have links isomorphic to $\Delta$. Before listing the definition of this cubical complex, we define what we mean by a cell complex in our setting.
@@
    By a cell complex, we mean a collection of subsets of the vertex set satisfying the formal properties of a polyhedral complex (Definition 5.1 on p. 127 of [38]) or the (polar) dual of one (analogous to polar duals of polytopes, defined in Part 2) or a modification of one by one of the following transformations:
    • A direct generalization of Tchebyshev subdivisions from Definition 2.1 on p. 573 of [19] (see Definition 1.1 and Remark 1.30 for intuitive idea and comments on generalizations)
    • Suppose that we have a proper coloring of the 1-skeleton of the given (dual of) a polyhedral complex and that each vertex is assigned a $\pm$ sign. A “signed unused color complex” keeps track of $\pm$ signs of vertices on unused colors of a face of the starting complex (Definition 1.22). Note that the number of colors used will be fixed as part of the initial conditions.
@@
    We will take the dual $A^{*}$ of a cell complex $A$ of dimension D1D-1 to mean the cell complex where a face of dimension $j$ corresponds to the collection of facets of $A$ that contain a particular (Dj1)(D-j-1)-dimensional face of $A$. This is an analogue of the polar dual of a polytope.
Remark 1.4.
The definitions above generalize directly to multicomplexes studied in [6]. We mention this since the identities of the form $h(\Delta)=f(\Gamma)$ used in our analysis (e.g. for flag Cohen–Macaulay simplicial complexes from [8]) also exist for multiplicomplexes.
We now give the definition of the cubical complex mentioned above.
@@
partially ordered by $0>1$ and 0>10>-1 extended componentwise. Here $I$ is the poset $\{0,1,-1\}$ with ordering $1<0$ and 1<0-1<0. Note that $\dim MT=\dim T+1$.
@@
(Comments on notation)
The way we will think about the notation above is that placing $1$ or 1-1 in the $i^{\text{th}}$ coordinate refers to intersecting with the hyperplane $x_{i}=1$ or $x_{i}=-1$ defining a facet of the $n$-cube $1,1]^{n}$ and that the coordinates labeled with a “0” are “free” in the sense that they are used to span a face of the cubical complex $MT$. In Rowlands’ thesis (see p. 4 of [31]), the “free” coordinates are labeled with a “*”, which may better convey this. As a consequence, the faces of $MT$ with 0’s in the same coordinates (i.e. coming from the same face of $T$) are parallel to each other.
Remark 1.7.
(Flag simplicial spheres to $CAT(0)$ complexes)
The mirroring/Danzer/power complex gives rise to a construction taking flag simplicial spheres to $CAT(0)$ complexes with repeated local structure given by $\Delta$ This “local structure” construction is motivated by a concrete construction using flag simplicial spheres $\Delta$ to build $CAT(0)$-complexes with repeated local structure given by $\Delta$ (Proposition 1.2.2 on p. 11 and Proposition 1.2.3 on p. 13 of [13], p. 21 of [12]).
@@
    Suppose that the faces $G_{1},\ldots,G_{m}\in MT$ have a nonempty intersection. For each $1\leq i\leq m$, let $F_{i}=\{(|a_{1}|,\ldots,|a_{n}|):(a_{1},\ldots,a_{n})\in G_{i}\}$ (i.e. subset labeled by 0’s corresponds to $F_{i}$). Then, we have that $G_{1}\cap\cdots\cap G_{m}$ has 0’s in the same slots as $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the “common values” from Part 1 for coordinates corresponding to elements of the complement $n]\setminus(F_{1}\cap\cdots\cap F_{m})$. In particular, this means that $F_{1}\cap\cdots\cap F_{m}\in M\Delta$ since the subset of $[n]$ labeled with 0’s is the face $F_{1}\cap\cdots\cap F_{m}\in\Delta$ and the rest of the coordinates are labeled with $\pm 1$.
@@
    • $\operatorname{st}_{M\Delta}(G)\cong\operatorname{st}_{\Delta}(|G|)$, where $|G|$ denotes the face of $\Delta$ obtained after taking the absolute value of each coordinate (i.e. setting the 1-1 coordinates of $G$ equal to 1).
    • The nonzero coordinates of elements of $\operatorname{st}_{M\Delta}(G)$ are determined by those of $G$ (which are indexed by $n]\setminus|G|$).
Returning to the original question of the interpretation of $P_{\Delta}(2\alpha+1)$, we use the following result connecting faces of a cell complex with its mirroring/Danzer/power complex.
@@
Suppose that we include the empty set in the face poset of a cell complex $T$ and that it is not included in the face poset of $MT$. If every (k1)(k-1)-dimensional cell of $T$ has $k$ vertices, we have that
@@
1.2.2 Colorings, Tchebyshev subdivisions, and intervals
The main aim of this section is to connect inverted Chebyshev expansions associated to gamma polynomials of flag simplicial spheres to cell complex structures. Our main tools are the compatibility between Tchebyshev subdivisions and $F$-polynomials from Section LABEL:tchebtypeAtoB and the $h$-vector itself being equal to the $f$-vector of an auxiliary simplicial complex in this setting (p. 470 of [8], Theorem 1 on p. 23 and Section 6.4 on p. 33 of [6]). One can think about the type $A$ to type $B$ transitions in terms of converting information about “Boolean structures” (e.g. face posets of simplicial complexes) into those about cubical complexes via repeated edge subdivisions. We start by considering flag simplicial spheres $\Delta$ where $\gamma(\Delta)=f(T)$ (e.g. barycentric subdivisions of simplicial spheres [27]). This will introduce transformations of cubical complexes compatible with those of the polynomials we study (e.g. translations) that describe the simplicial to cubical complex transitions in a more concrete way.
Note that having $h(\Delta)=f(\Gamma)$ is a necessary condition in order for this occur (Proposition 6.2 on p. 1377 of [27]). When $\Delta$ is a flag sphere, this equality holds and $\Gamma$ can be taken to be a balanced simplicial complex. Combining this with the compatibility between Tchebyshev transformations and $F$-polynomials, we obtain our main formula connecting the gamma polynomial to $F$-polynomials of an auxiliary cell complex coming from signed colorings of $\Gamma$ (Corollary 1.33). The $F$-polynomial is the translation by -1 of the generating function for the number of faces of a cubical complex whose local structure is this determined by this auxiliary cell complex. The motivation comes a certain construction that converts flag simplicial spheres $\Delta$ into $CAT(0)$-complexes whose local structure is given by $\Delta$ (Proposition 1.2.2 on p. 11 and Proposition 1.2.3 on p. 13 of [13], p. 21 of [12], p. 302 of [5]).
Finally, we use this formula to show new positivity properties involving the reciprocal polynomial of the gamma polynomial that hold for arbitrary flag spheres and that the $f$-polynomial of this cubical complex “formally” behaves like a cubical version of a barycentric subdivision (where the word “formally” can be removed when $\gamma(\Delta)=f(T)$) (Corollary 1.35). Note that the this comes from a translation, which is also related to transformations between $F$-polynomials and $f$-polynomials. This is followed by some further context involving earlier work studying connections between $CAT(0)$ cubical complexes and flag simplicial complexes.
Now suppose that $\gamma(\Delta)=f(T)$ for some simplicial complex $T$ of dimension $\leq\frac{d}{2}-1$. By Theorem 0.1, we have that
@@ 1.2.2 Colorings, Tchebyshev subdivisions, and intervals
The first cubical barycentric subdivision of a simplicial complex $S$ of dimension D1D-1 is the cubical complex $\mathcal{C}(S)$ defined as follows:
@@
    $\sigma,\tau]\coloneq\{\lambda\in S:\sigma\subset\lambda\subset\tau\}.$
@@
(Comparison with the formula for $w^{k}]g(2w)$ given if $\gamma(\Delta)=f(T)$)
@@
    Comparing this to the formula for $w^{k}]g(2w)$ obtained when $\Gamma(\Delta)=f(T)$, we see that it uses $\binom{j}{k}$ instead of $\binom{j+1}{k}$ and has an input that might not be a simplicial complex. Note that the $k$-dimensional faces of $MA$ come from (k1)(k-1)-dimensional faces of $A$ (see Definition 1.5).
@@
    The cubical barycentric subdivision assumes that both endpoints of the interval considered are contained in $S$. If we drop the condition that $S$ is a simplicial complex, we’re simply considering endpoints $\sigma,\tau]$ where the upper endpoint is a $j$-dimensional face of $S$ and $\tau$ is some subset of $\sigma$ (which may or may not be a face of $S$).
@@
Given a poset $P$, its barycentric cover KPKP is the poset with elements the order relations of $P$, partially ordered by inclusion (i.e. $(u\leq v)\leq(x\leq y)$ (see Definition 1.17) if and only if $x\leq u\leq v\leq y$). Note that $KP=K^{\text{op}}$ (see p. 298 of [5]) and $K(P\times Q)=K(P)\times K(Q)$.
@@
This explains the “cubical” part and the induced triangulation of cubes in $KP$ induced by the barycentric subdivision give the cubical barycentri subdivsion name (see top of p. 304 of [5]). More information on structural properties of $KP$ are listed below.
@@
An interval $[u,v]$ in a partially ordered set $P$ is the set of all elements $w\in P$ satisfying $u\leq v\leq w$. For a finite partially ordered set $P$, we define the poset $I(P)$ of the nonempty intervals of $P$ as the set of all intervals $u,v]\subset P$, ordered by inclusion.
@@
For each $u\in P$, we identify the vertex $u,u]\in\Delta(I(P))$ with the vertex $u\in\Delta(P)$ and for each nonsingleton interval $u,v]\in I(P)$, we identify the vertex $u,v]\in\Delta(I(P))$ with the midpoint of the edge $\{[u,u],[v,v]\}$. We number the midpoints $u_{1},v_{1}],[u_{2},v_{2}],\ldots$ in such a way that $i<j$ holds whenever the interval $u_{i},v_{i}]$ contains the interval $u_{j},v_{j}]$.
@@
[…]
@@
We take a page from p. 13 of Rowlands’ paper [32], which shows that $\widetilde{f}(\Box_{P},t)=\widetilde{f}(\Delta_{P},t+1)$. Note that any flag complex is the crossing complex $\Delta_{P}$ of a PIPPIP $P$ (Lemma 3.3 on p. 11 of [32]) and any rooted $CAT(0)$ cubical complex is of the form $\Box_{P}$ for some $P$ (Theorem 2.9 on p. 9 of [32]). The proof uses a bijection between dimension $i$ elements of $\Box_{P}$ and pairs $(A,M)$, where $A$ is the set of consistent antichains of $P$ and $M\subset A$ is a subset with $|M|=i$. The identity follows after making a change in the order of summation.
@@
$w^{k}]\widetilde{f}\left(T(A),\frac{w-1}{2}\right)=\sum_{j=k}^{\frac{d}{2}}\widetilde{f}_{j}(T^{*})2^{j}\binom{j}{k}$,
where $A$ is cell complex related to unused colors of faces in a balanced simplicial complex with an assigned coloring. Note that the last expression makes sense since a (k1)(k-1)-face of $T(A)$ has $k$ vertices (see Definition 1.29).
@@
Let $\Gamma$ be a (D1)(D-1)-dimensional balanced simplicial complex. Fix a $D$-coloring of the $1$-skeleton of $\Gamma$. We define the signed unused color complex $\mathcal{D}(\Gamma)$ of $\Gamma$ to be the cell complex below.
@@
  • The (k1)(k-1)-faces of the form $B\subset Q\subset C_{F}$, where $C_{F}$ is the set of colors in $[D]$ not used by $F$, $Q\subset C_{F}$ is a $k$-element subset, and $B\subset Q$ is an arbitrary subset. This is like considering “signed” $k$-element subsets of $C_{F}$. Note that $C_{F}\neq\emptyset$ since $Q$ is nonempty for the cases we will consider.
@@
Let $\Gamma$ be a balanced simplicial complex of dimension D1D-1. A collection of vertices $v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\in V(\mathcal{D}(\Gamma))$ with each $v_{i}$ in a triple with a face $F_{i}\in\Gamma$ and sign $\varepsilon_{i}=\pm 1$ forms a face of $\mathcal{D}(\Gamma)$ using the union of set of unused colors $C_{i}$ of the $F_{i}$ if and only if one of the following conditions (in the notation of Definition 1.22) holds:
@@
The second condition is nontrivial. For example, suppose that $\Gamma$ is the boundary of the octahedron/2-cross polytope (meaning $D=3$). This is a balanced simplicial complex. Color the vertices $\pm e_{i}$ with $i$ for $1\leq i\leq 3$ by $i$ to get a 3-coloring of its 1-skeleton. Consider the vertex $F_{1}=\{-e_{2}\}$ and the edge $F_{2}=\{e_{2},e_{3}\}$. Then, we have that $C_{F_{1}}=\{1,3\}$ and $C_{F_{2}}=\{1\}$. Since $F_{1}\cap F_{2}=\emptyset$ and $C_{F_{1}}\cup C_{F_{2}}=\{1,3\}$, the union of any two vertices of $\mathcal{D}(\Gamma)$ associated to these faces cannot form a face of $\mathcal{D}(\Gamma)$.
@@
Each vertex is of the form $B_{i}\subset\{q_{i}\}\subset C_{F_{i}}$, where $C_{F_{i}}$ is the set of unused colors of $F_{i}\in\Gamma$, $q_{i}\in C_{F_{i}}$, and $B_{i}=\{q_{i}\}$ or $\emptyset$. Using the inclusion relation in Definition 1.22, we have that the union $\{v_{1}^{\varepsilon_{1}},\ldots,v_{k}^{\varepsilon_{k}}\}$ is given by $B_{1}\cup\cdots\cup B_{k}\subset\{q_{1},\ldots,q_{k}\}\subset(C_{1}\cup\cdots\cup C_{k})_{F_{1}\cap\cdots\cap F_{k}}$. If $F_{1}\cap\cdots\cap F_{k}\neq\emptyset$, this forms a face of $\mathcal{D}(\Gamma)$ since $(P\cap Q)^{c}=P^{c}\cup Q^{c}$ for $P,Q\subset[D]$. Now suppose that $F_{1}\cap\cdots\cap F_{k}=\emptyset$. As mentioned in Definition 1.22, we have that $C_{\emptyset}=[D]$. This is only possible if $C_{F_{1}}\cup\cdots\cup C_{F_{k}}=[D]$.
@@
$\alpha^{k}]\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}=f_{k-1}(\mathcal{D}(\Gamma))$
for $k\geq 1$ and
@@
for a cell complex $A$ of dimension D1D-1.
@@
Since $\Delta$ is a flag simplicial homology sphere, it is also a flag Cohen–Macaulay simplicial complex (e.g. see p. 21 of [4]). Then, a combining a result of Caviglia–Constantinescu–Varbaro (Corollary 2.3 on p. 474 of [8]) with that of Björner–Frankl–Stanley ($a=(1,\ldots,1)$ case of Theorem 1 on p. 23 and Corollary from Section 6.4 on p. 33 of [6]) implies that $h(\Delta)$ is the $f$-vector of a balanced $\left(\frac{d}{2}-1\right)$-dimensional simplicial complex $\Gamma$. Note that the indexing convention for $f$-vectors on p. 24 and Section 6.4 on p. 33 of [6] differs from the usual one for the dimensions of faces of simplicial complexes and is closer to those of cubical complexes if empty sets were to be included. We will use the usual convention for $f$-vectors of simplicial complexes by indexing the faces by dimensions instead of number of vertices. From now on, we will write $\Gamma$ to mean that $\left(\frac{d}{2}-1\right)$-skeleton of this simplicial complex.
@@
$\frac{P_{\Delta}(2\alpha+1)+h_{\frac{d}{2}}(\Delta)}{2}$ $=\frac{P_{\Delta}(2\alpha+1)+f_{\frac{d}{2}-1}(\Gamma)}{2}$
$=f_{\frac{d}{2}-1}(\Gamma)+\sum_{j=1}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)(2\alpha+1)^{j}$
$=\sum_{j=0}^{\frac{d}{2}}f_{\frac{d}{2}-j-1}(\Gamma)(2\alpha+1)^{j}$